OpenStudy (anonymous):

Solve for θ in the equation tan θ = -4 when 0º < θ < 180º. Round your answer to the nearest tenth of a degree.

OpenStudy (anonymous):

cotθ = 2 1/tanθ = 2 tanθ = 1/2 θ = arctan(1/2), arctan(1/2) + π θ = 0.463647609, 3.605240263 Source yahoo answers

OpenStudy (anonymous):

Nice @Rohangrr

OpenStudy (anonymous):

i have to find my answer in degrees like my choices are 99.5 104.0 106.4 100.4

OpenStudy (anonymous):

iam stuck on this one

OpenStudy (anonymous):

its then 104

OpenStudy (anonymous):

thanks i dont have that graphing calucator

OpenStudy (anonymous):

Nice rohangrr